Question: Annual demand for an item is 15,600. Weekly demand is 300 units. The cost of placing an order is $100 and the time from ordering to receipt is two weeks. Annual inventory carrying cost is $1 per unit. Find the economic order quantity and reorder point for this item.

Answer: Q* = 1,766.35 ~ 1,766; r = DL = 600.

Professor gave answer, but I would like to see the steps on how to get the answer.
Answer #1

1. DEMAND = 15600
ORDERING COST = 100
HOLDING COST = 1


EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST) = SQRT(2 * 15600 * 100 / 1) = 1766.35

WEEKLY DEMAND = 300

REORDER POINT = DEMAND DURING LEAD TIME
= 300 * 2 = 600